OpenStudy (anonymous):

Statistics: A batch consists of 12 defective coils and 88 good ones. Find the probability of getting two good coils when two coils are randomly selected if the first selection is replaced before the second is made. A) 0.7733 B) 0.176 C) 0.0144 D) 0.7744

jimthompson5910 (jim_thompson5910):

Hint: P(2 good coils) = P(Good coil and Good coil) P(2 good coils) = P(Good coil) * P(Good coil)

OpenStudy (anonymous):

i dont get it...

jimthompson5910 (jim_thompson5910):

What's P(good coil)

OpenStudy (anonymous):

88/100?

jimthompson5910 (jim_thompson5910):

so... P(2 good coils) = P(Good coil and Good coil) P(2 good coils) = P(Good coil) * P(Good coil) P(2 good coils) = (88/100) * (88/100) P(2 good coils) = ???

OpenStudy (anonymous):

ah ok, i get it

jimthompson5910 (jim_thompson5910):

ok what answer did you get

OpenStudy (anonymous):

0.7744

jimthompson5910 (jim_thompson5910):

yep you got it